OpenStudy (anonymous):

PLz help asap gold medal will be awarded.find f(5) for this sequence. f(1)=2and f(2)=5,f(n)= f(1) +f(2)+f(n-1), for n>2 f(5)=....

OpenStudy (anonymous):

Use the formula, and let n = 3, so f(3) = f(1) + f(2) + f(2) = 2 + 5 + 5 = 12 Now, let's put in n = 4, we get f(4) = F(1) + f(2) + f(3) = 2 + 5 + 12 = 19 Now, we go to n = 5 (which we want in this problem): f(5) = f(1) + f(2) + f(4) = 2 + 5 + 19 = 26 Hope this helps.

OpenStudy (anonymous):

um yes it does thanks

OpenStudy (anonymous):

Welcome. If there's anything that you need clarification, feel free to write to me.

OpenStudy (anonymous):

ok i will thx

OpenStudy (anonymous):

Short and sweet....we were given f(1) and f(2), and we waanted to find f(5). So we first find f(3), then f(4), then we can finally find f(5). and we use the rule to find f(3), f(4), and f(5).